OpenStudy (anonymous):

integrate 1/cscx - 1 dx

myininaya (myininaya):

\[\int\limits_{}^{}(\frac{1}{cscx}-1) dx\] correct?

myininaya (myininaya):

\[\int\limits_{}^{}(sinx-1)dx=\int\limits_{}^{}sinx dx -\int\limits_{}^{}1dx+C=-cosx-x+C\]

OpenStudy (anonymous):

i thought i was doing it wrong...

myininaya (myininaya):

so you were doing it right?

OpenStudy (anonymous):

yes... i still fell insecure of what to do sometimes,, and i have a test in few hours

myininaya (myininaya):

why did you feel insecure about this problem?

OpenStudy (anonymous):

its not only about this problem... is what to do when i see different problems.. i think,, maybe i am wrong,,, and i delete everything and start another mess... like now... and ive been practice the whole week,, but i dont know,,,

myininaya (myininaya):

it all takes practice trail and error

OpenStudy (anonymous):

when you can do the ones in your calc I book in your head you can be sure you know what your doing